Problem Case # A. Celling Fun The next nine questions pertain to the situation described below. 110] A ceiling fan consists of a small cylindrical disk with 5 thin rods coming from the center. The disk has mass m 3.3 kg and adius R- 0.22 m The rods each have mass m- 1.2 kg and length L-0.78 m 1 What is the moment of inertia of each rod about the axis o rotation? 2) What is the moment of inertis of the disk about the axis of rotation? What is the moment of inertia of the whole ceiling fan?
Answer #1

1) Ir = mr L^2/3 = 1.2*0.78*0.78/3 = 0.24 Kg.m2

2) Id = md R^2/2 = 3.3*0.22*0.22/2 = 0.08 Kg.m2

3)I = 0.24 + 0.08 = 0.32 Kg.m2

4) theta = 16 * (2*3.14) = 100.48 radians

Theta = 0.5 ? t^2

Alpha = 2 theta/t^2 = 2 * 100.48/(3.5*3.5) = 16. 40 rad/s2

5) Omega = alpha x t = 16. 40* 3.5 = 57. 42 rad/s

According to Chegg guidelines I can answer these many sub parts in a question, so please post other parts in a separate question.